# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 5.2 Datasets
Datasets for Continual Multimodal Large Language Models. Following LLaVA [167], many MLLMs adopt the pattern of instruction tuning to enable assessing alignment with human intention and knowledge preservation for reasoning. Thus, traditional tasks like image classification can be transformed to VQA tasks to evaluate the ability of MLLMs, which are otherwise challenging to assess using conventional methods. Several benchmarks have been proposed to evaluate the CL method for MLLMs. MCIT [100] proposes the first continual instruction tuning benchmarks, Benchmark1 and Benchmark2. The difference between benchmark1 and benchmark2 is that benchmark2 includes Multi-task Joint Instruction Tuning, which aims to explore whether multi-task joint instruction tuning improves the model’s continual learning ability. [349] proposes EMT, the first classification evaluation framework to investigate catastrophic forgetting in MLLMs. [41] presents a comprehensive benchmark CoIN, spanning 8 task categories and evaluating MLLMs from two perspectives: Instruction Following and General Knowledge, which assess the alignment with human intention and knowledge preserved for reasoning, respectively. [362] constructs two datasets, UPMC-Food101-CMML and MM-IMDb-CMML to benchmark the novel CMML task, which means the data of certain modalities is missing during continual fine-tuning. UPMC-Food101-CMM contains 101 food categories and 61,142 training, 6,846 validation, and 22,716 test image-text pairs. MM-IMDb-CMML is a multi-label classification dataset across 27 distinct movie genres, consisting of 15,552 training, 2,608 validation and 7,799 test image-text pairs.
6 Discussion
@@
6.1 Intriguing Properties Emergent in Continual LLMs
@@
Beyond the well-established resilience of pre-trained large language models (LLMs) against catastrophic forgetting compared to downstream-specific models [133, 274, 183, 365, 195], there is a notable lack of exploration into other intriguing properties of LLMs when trained continually. While investigations into the emergent capabilities of continuously trained LLMs have attracted attentions from the community to a certain extent, they remain relatively limited. For instance, in [332], it is observed that when fine-tuned sequentially and cyclically on a series of documents, large models exhibit a phenomenon known as “anticipatory recovering”. This refers to the LLMs’ ability to recover forgotten information on documents even before encountering them again. This suggests that LLMs may possess the capability of sequential memorization, which could pave the way for research into memory replay and more complex structured learning environments as model parameters scale up.
6.2 Conventional Types of Incremental Learning
@@
As mentioned in Section 2.2, three types of incremental learning are prevalent [282]. Among them, class-incremental learning (CIL) has historically attracted significant attention from the community [239, 316]. However, in the context of continually pre-training and adapting large language models (LLMs), we observe a decreased interest in CIL but an increased focus on task-incremental learning (TIL) and domain-incremental learning (DIL). Given that language models are inherently designed for content generation and are pre-trained with the pretext generative task of next-word prediction, it is natural to emphasize the patterns of generative tasks and integrate the traditional CIL paradigm into the broader framework of language modeling, discarding the incremental classification head [258, 57, 33]. For instance, in Vocabulary-Aware Label Generation (VAG), CIL is redefined as the task of continual label generation. This approach utilizes a pre-trained encoder-decoder language model to generate class labels [258]. Meanwhile, in the Generative Multi-modal Model (GMM) for CIL [33], image patches and prompts are concatenated and fed into the language model to generate classification results.
@@
However, the declining attention to the conventional CIL paradigm does not suggest that these techniques are not impactful in the field of continual learning for LLMs. Nonetheless, many current research endeavors unwittingly employ such techniques, indicating their widespread adoption in various applications. For example, techniques such as vocabulary expansion [6, 55] can be seen as an extension of expanding the classification head in CIL. These CIL techniques can be further integrated into systems like Lifelong-MoE [46], where adding a new expert to the transformer blocks requires updating the gating function to include the routing of the newly added expert. The EProj framework, described in [100], employs a similar architecture, incorporating linear projectors for new domains alongside a selector module trained to route these projectors. Since the aforementioned sub-modules operate on the principles of CIL, previously validated techniques can be directly applied.
@@
The importance of domain-incremental learning (DIL) is self-evident, given the shared task definition and input-output format in continual pre-training (CPT) and domain-adaptive pre-training (DAP). As dynamically expanding token vocabularies can pose additional challenges, it is natural to focus on understanding distributional shifts within the input corpus while keeping the vocabulary fixed. On another front, task-incremental learning (TIL) attracts significant interest due to its potential for personalizing LLM services. For instance, users may desire options for selecting domain-specific experts, thereby making task IDs available throughout inference time [112, 309]. Additionally, TIL plays a crucial role in instruction tuning, where instructions can be seen as natural-language-encoded task information [255, 113, 205, 103, 337, 291, 363, 294]. It is worth noting that the boundary between TIL and DIL becomes somewhat blurred in continual instruction tuning. Language models demonstrate the capability to infer domain information for unseen instructions, suggesting a convergence of TIL and DIL in certain contexts.
6.3 Roles of Memory in Continual LLMs
@@
Previous continual learning research, drawing inspiration from human learning patterns, primarily emphasizes the storage efficiency of past data. The setting of continual learning with limited memory size has garnered significant attention from the community. However, this focus may no longer hold true in the context of continual LLMs. In the direction of relaxing memory constraints, institutions with access to training data may opt to retain full access without restricting memory size, given that the cost of memory storage is more than affordable. In such scenarios, as highlighted in [283], the challenge shifts from storage efficiency to computational efficiency. To achieve continual learning goals, models must efficiently adapt to new data (efficient adaptation) and select key experiences for replay (efficient replay) [323, 126]. Therefore, it is essential to reassess the existing memory constraint and prioritize optimizing computational efficiency for continual learning of LLMs by restricting the number of updates and the number of FLOPs [227, 298].
@@
On the other end of the spectrum, studies with tightened memory constraints remain vital in modern continual learning of LLMs. As shown in Fig. 1, upstream suppliers of LLMs typically do not provide training data with the released model weights. Consequently, consumers must adapt these models to downstream data without access to the actual replay data. Various rehearsal-free continual strategies are applied in this scenario, such as collecting data examples from alternate sources [245, 52, 311, 10], leveraging the generative capabilities of LLMs to produce pseudo-examples for replay [230], and implementing regularization techniques in the parameter space [134, 243]. Continual learning under the strict memory constraint is also driven by data privacy concerns, where preserving data on the server side is prohibited. In these scenarios, researchers must rely on online continual learning methods [31, 189, 228], where data examples are only utilized for training as they arrive in a stream, and numerous efforts are already underway to develop LLMs capable of operating under these constraints [329, 297, 25].
6.4 Prospective Directions
@@
Theories of Continual LLMs. It is widely recognized that the continual learning community tends to prioritize empirical research over theoretical exploration. Nevertheless, there are efforts to establish theoretical foundations for CL. In [288], the authors utilize second-order Taylor expansions around optimal parameters to derive an inter-task generalization error bound based on the maximum eigenvalue and $l_{2}$-norm of parameter differences. Another line of approaches leverages task/domain discrepancies to construct a multi-task generalization bound. For instance, Unified Domain Incremental Learning (UDIL) in [260] proposes upper bounds for intra-domain and cross-domain distillation losses, unifying various replay-based DIL techniques under a single adaptive generalization bound. However, applying these existing theories directly to continual LLMs can be imprudent, given their pre-trained, large-scale nature. Consequently, there is a notable gap in research focusing on continually learning LLMs with robust theoretical guarantees and understanding the forgetting behaviors of LLMs from a theoretical perspective.
@@
Efficient Replay for Knowledge Retention for Continual LLMs. Computational resources for training large-scale LLMs are often limited. While the storage budget can theoretically be infinite (Section 6.3), replaying past experiences without specific design can lead to inefficient updates in current domain learning, resulting in slow convergence. Beyond sparse replay solutions that control data mixture ratios [166, 245, 331], there is ongoing exploration of efficient replay for continual LLMs. For example, KPIG [103] enhances replay efficiency by calculating Key-part Information Gain (KPIG) on masked segments, enabling the dynamic selection of replay data. In [126], a pioneering effort introduces a forgetting forecasting mechanism based on output changes during adaptation, later used for selective replay in continual model refinement (CMR). It has been verified in this work that filtering replay samples based on their tendency to forget significantly improves knowledge retention rates for continual LLMs. However, more sophisticated and accurate data mixing strategies and efficient replay sample selection mechanisms are much needed, e.g., a dynamic data mixing ratio throughout the training process. Hence we mark this practical direction of efficient replay for LLMs a significant research focus in the future.
@@
Continual LLMs with Controllable Memory. The long-term memory inherent in the whole set of parameters of LLMs often lacks interpretability and explicit manipulability, which is crucial in certain application areas. For instance, consider a scenario where a supplier collects data from customers under their consent and continually utilizes this data to update LLMs. However, if some users later revoke their consent, the knowledge acquired by the trained model from that portion of data must also be revoked. With a continually pre-trained large-scale LLM, the only solution is to roll back to a previous model version predating the inclusion of these users’ data and retrain the model from that point onward. This example of “machine unlearning” [26, 210] vividly illustrates the benefits of equipping LLMs with an external, controllable memory. As part of continual model refinement (CMR), memory systems for continual learning have been explored in several studies. Larimar [58] suggests integrating the Kanerva Machine [317] as an episodic memory for multi-fact model editing. This memory system supports basic operations like writing, reading, and generating, as well as advanced operations such as sequential writing and forgetting. It enables one-shot knowledge updates without costly retraining or fine-tuning. Additionally, other memory systems like Hopfield Networks [238, 226] hold promise for future investigation.
@@
Continual LLMs with Custom Preferences. Customizing user preferences is critical for LLMs, especially in service-oriented contexts. Users often require different trade-offs between domain expertise, ethics, values, or tones of expression. Efficiently building customized LLMs for individual users and offering flexible adjustment options is a challenging task. Early attempts in this direction include Imprecise Bayesian Continual Learning (IBCL), which, under certain assumptions, guarantees the generation of Pareto-optimal models based on user preferences by combining two model posteriors in the parameter space [178]. While empirical validation is limited in scale, this approach paves the way for future research in this area.
7 Conclusion

### v2

@@ 5.2. Datasets
Datasets for Continual Multimodal Large Language Models. Following LLaVA (liu2023visual, ), many MLLMs adopt the pattern of instruction tuning to enable assessing alignment with human intention and knowledge preservation for reasoning. Thus, traditional tasks like image classification can be transformed to VQA tasks to evaluate the ability of MLLMs, which are otherwise challenging to assess using conventional methods. Several benchmarks have been proposed to evaluate the CL method for MLLMs. MCIT (he2023continual, ) proposes the first continual instruction tuning benchmarks, Benchmark1 and Benchmark2. The difference between benchmark1 and benchmark2 is that benchmark2 includes Multi-task Joint Instruction Tuning, which aims to explore whether multi-task joint instruction tuning improves the model’s continual learning ability. (zhai2023investigating, ) proposes EMT, the first classification evaluation framework to investigate catastrophic forgetting in MLLMs. (chen2024coin, ) presents a comprehensive benchmark CoIN, spanning 8 task categories and evaluating MLLMs from two perspectives: Instruction Following and General Knowledge, which assess the alignment with human intention and knowledge preserved for reasoning, respectively. (zhao2024reconstruct, ) constructs two datasets, UPMC-Food101-CMML and MM-IMDb-CMML to benchmark the novel CMML task, which means the data of certain modalities is missing during continual fine-tuning. UPMC-Food101-CMM contains 101 food categories and 61,142 training, 6,846 validation, and 22,716 test image-text pairs. MM-IMDb-CMML is a multi-label classification dataset across 27 distinct movie genres, consisting of 15,552 training, 2,608 validation and 7,799 test image-text pairs.
6. Discussion
@@
6.1. Intriguing Properties Emergent in Continual LLMs
@@
Beyond the well-established resilience of pre-trained large language models (LLMs) against catastrophic forgetting compared to downstream-specific models (ke2021achieve, ; tao2022can, ; luo2023investigating, ; zheng2023learn, ; mehta2023empirical, ), there is a notable lack of exploration into other intriguing properties of LLMs when trained continually. While investigations into the emergent capabilities of continuously trained LLMs have attracted attentions from the community to a certain extent, they remain relatively limited. For instance, in (yang2024reawakening, ), it is observed that when fine-tuned sequentially and cyclically on a series of documents, large models exhibit a phenomenon known as “anticipatory recovering”. This refers to the LLMs’ ability to recover forgotten information on documents even before encountering them again. This suggests that LLMs may possess the capability of sequential memorization, which could pave the way for research into memory replay and more complex structured learning environments as model parameters scale up.
6.2. Conventional Types of Incremental Learning
@@
As mentioned in Section 2.2, three types of incremental learning are prevalent (van2022three, ). Among them, class-incremental learning (CIL) has historically attracted significant attention from the community (rebuffi2017icarl, ; wu2019large, ). However, in the context of continually pre-training and adapting large language models (LLMs), we observe a decreased interest in CIL but an increased focus on task-incremental learning (TIL) and domain-incremental learning (DIL). Given that language models are inherently designed for content generation and are pre-trained with the pretext generative task of next-word prediction, it is natural to emphasize the patterns of generative tasks and integrate the traditional CIL paradigm into the broader framework of language modeling, discarding the incremental classification head (shao2023class, ; dalessandro2023multimodal, ; cao2024generative, ). For instance, in Vocabulary-Aware Label Generation (VAG), CIL is redefined as the task of continual label generation. This approach utilizes a pre-trained encoder-decoder language model to generate class labels (shao2023class, ). Meanwhile, in the Generative Multi-modal Model (GMM) for CIL (cao2024generative, ), image patches and prompts are concatenated and fed into the language model to generate classification results.
@@
However, the declining attention to the conventional CIL paradigm does not suggest that these techniques are not impactful in the field of continual learning for LLMs. Nonetheless, many current research endeavors unwittingly employ such techniques, indicating their widespread adoption in various applications. For example, techniques such as vocabulary expansion (amba2021dynamic, ; cossu2022continual, ) can be seen as an extension of expanding the classification head in CIL. These CIL techniques can be further integrated into systems like Lifelong-MoE (chen2023lifelong, ), where adding a new expert to the transformer blocks requires updating the gating function to include the routing of the newly added expert. The EProj framework, described in (he2023continual, ), employs a similar architecture, incorporating linear projectors for new domains alongside a selector module trained to route these projectors. Since the aforementioned sub-modules operate on the principles of CIL, previously validated techniques can be directly applied.
@@
The importance of domain-incremental learning (DIL) is self-evident, given the shared task definition and input-output format in continual pre-training (CPT) and domain-adaptive pre-training (DAP). As dynamically expanding token vocabularies can pose additional challenges, it is natural to focus on understanding distributional shifts within the input corpus while keeping the vocabulary fixed. On another front, task-incremental learning (TIL) attracts significant interest due to its potential for personalizing LLM services. For instance, users may desire options for selecting domain-specific experts, thereby making task IDs available throughout inference time (huang2023lorahub, ; wistuba2023continual, ). Additionally, TIL plays a crucial role in instruction tuning, where instructions can be seen as natural-language-encoded task information (scialom2022fine, ; huang2024mitigating, ; mok2023large, ; he2024dont, ; yin2022contintin, ; wang2023orthogonal, ; zhao2024sapt, ; wang2024inscl, ). It is worth noting that the boundary between TIL and DIL becomes somewhat blurred in continual instruction tuning. Language models demonstrate the capability to infer domain information for unseen instructions, suggesting a convergence of TIL and DIL in certain contexts.
6.3. Roles of Memory in Continual LLMs
@@
Previous continual learning research, drawing inspiration from human learning patterns, primarily emphasizes the storage efficiency of past data. The setting of continual learning with limited memory size has garnered significant attention from the community. However, this focus may no longer hold true in the context of continual LLMs. In the direction of relaxing memory constraints, institutions with access to training data may opt to retain full access without restricting memory size, given that the cost of memory storage is more than affordable. In such scenarios, as highlighted in (verwimp2024continual, ), the challenge shifts from storage efficiency to computational efficiency. To achieve continual learning goals, models must efficiently adapt to new data (efficient adaptation) and select key experiences for replay (efficient replay) (xie2023efficient, ; jin2024model, ). Therefore, it is essential to reassess the existing memory constraint and prioritize optimizing computational efficiency for continual learning of LLMs by restricting the number of updates and the number of FLOPs (prabhu2023computationally, ; wang2022sparcl, ).
@@
On the other end of the spectrum, studies with tightened memory constraints remain vital in modern continual learning of LLMs. As shown in Fig. 1, upstream suppliers of LLMs typically do not provide training data with the released model weights. Consequently, consumers must adapt these models to downstream data without access to the actual replay data. Various rehearsal-free continual strategies are applied in this scenario, such as collecting data examples from alternate sources (rozière2024code, ; colombo2024saullm7b, ; wu2023pmc, ; Azerbayev2023LLEMMA, ), leveraging the generative capabilities of LLMs to produce pseudo-examples for replay (qin2021lfpt5, ), and implementing regularization techniques in the parameter space (ke2022continual-pre, ; rongali2021continual, ). Continual learning under the strict memory constraint is also driven by data privacy concerns, where preserving data on the server side is prohibited. In these scenarios, researchers must rely on online continual learning methods (cai2021online, ; mai2022online, ; prabhu2023online, ), where data examples are only utilized for training as they arrive in a stream, and numerous efforts are already underway to develop LLMs capable of operating under these constraints (yang2022continual, ; wang2022online, ; bornschein2024transformers, ).
6.4. Prospective Directions
@@
Theories of Continual LLMs. It is widely recognized that the continual learning community tends to prioritize empirical research over theoretical exploration. Nevertheless, there are efforts to establish theoretical foundations for CL. In (wang2024comprehensive, ), the authors utilize second-order Taylor expansions around optimal parameters to derive an inter-task generalization error bound based on the maximum eigenvalue and $l_{2}$-norm of parameter differences. Another line of approaches leverages task/domain discrepancies to construct a multi-task generalization bound. For instance, Unified Domain Incremental Learning (UDIL) in (shi2024unified, ) proposes upper bounds for intra-domain and cross-domain distillation losses, unifying various replay-based DIL techniques under a single adaptive generalization bound. However, applying these existing theories directly to continual LLMs can be imprudent, given their pre-trained, large-scale nature. Consequently, there is a notable gap in research focusing on continually learning LLMs with robust theoretical guarantees and understanding the forgetting behaviors of LLMs from a theoretical perspective.
@@
Efficient Replay for Knowledge Retention for Continual LLMs. Computational resources for training large-scale LLMs are often limited. While the storage budget can theoretically be infinite (Section 6.3), replaying past experiences without specific design can lead to inefficient updates in current domain learning, resulting in slow convergence. Beyond sparse replay solutions that control data mixture ratios (lin2023geogalactica, ; rozière2024code, ; Yang2023PLLaMa, ), there is ongoing exploration of efficient replay for continual LLMs. For example, KPIG (he2024dont, ) enhances replay efficiency by calculating Key-part Information Gain (KPIG) on masked segments, enabling the dynamic selection of replay data. In (jin2024model, ), a pioneering effort introduces a forgetting forecasting mechanism based on output changes during adaptation, later used for selective replay in continual model refinement (CMR). It has been verified in this work that filtering replay samples based on their tendency to forget significantly improves knowledge retention rates for continual LLMs. However, more sophisticated and accurate data mixing strategies and efficient replay sample selection mechanisms are much needed, e.g., a dynamic data mixing ratio throughout the training process. Hence we mark this practical direction of efficient replay for LLMs a significant research focus in the future.
@@
Continual LLMs with Controllable Memory. The long-term memory inherent in the whole set of parameters of LLMs often lacks interpretability and explicit manipulability, which is crucial in certain application areas. For instance, consider a scenario where a supplier collects data from customers under their consent and continually utilizes this data to update LLMs. However, if some users later revoke their consent, the knowledge acquired by the trained model from that portion of data must also be revoked. With a continually pre-trained large-scale LLM, the only solution is to roll back to a previous model version predating the inclusion of these users’ data and retrain the model from that point onward. This example of “machine unlearning” (bourtoule2020machine, ; nguyen2022survey, ) vividly illustrates the benefits of equipping LLMs with an external, controllable memory. As part of continual model refinement (CMR), memory systems for continual learning have been explored in several studies. Larimar (das2024larimar, ) suggests integrating the Kanerva Machine (wu2018kanerva, ) as an episodic memory for multi-fact model editing. This memory system supports basic operations like writing, reading, and generating, as well as advanced operations such as sequential writing and forgetting. It enables one-shot knowledge updates without costly retraining or fine-tuning. Additionally, other memory systems like Hopfield Networks (ramsauer2021hopfield, ; pourcel2022online, ) hold promise for future investigation.
@@
Continual LLMs with Custom Preferences. Customizing user preferences is critical for LLMs, especially in service-oriented contexts. Users often require different trade-offs between domain expertise, ethics, values, or tones of expression. Efficiently building customized LLMs for individual users and offering flexible adjustment options is a challenging task. Early attempts in this direction include Imprecise Bayesian Continual Learning (IBCL), which, under certain assumptions, guarantees the generation of Pareto-optimal models based on user preferences by combining two model posteriors in the parameter space (lu2023ibcl, ). While empirical validation is limited in scale, this approach paves the way for future research in this area.
7. Conclusion